OpenStudy (anonymous):

It was 41 degress F at 8:00 am and 69 degrees F at 4:00 pm If the temperature rose at a constant rate what was the temperature at 1:00 pm?

OpenStudy (anonymous):

Convert 4pm to 16am and 1pm to 13am. m= (69-41)/(16-8) = 28/8 = 7/2 (x-41)/(13-8) = 7/2 2(x-41) = 7(5) 2x - 82 = 35 2x = 117 x = 58.5 degrees F

OpenStudy (anonymous):

I don't understand the conversion of the time and where that comes from

OpenStudy (shane_b):

(Fix..though it was at 2pm!) The temperature went up 28 degrees in 8 hours changing at a constant rate...so the ratio is: \[\frac{28}{8}=\frac{x}{5}\] Solve for x to get a +17.5 degree increase. 41+17.5 = 58.5F.

OpenStudy (shane_b):

You don't need to bother with the time conversion if you just think about how many hours are between 8am and 4pm, and 8am and 1pm.

OpenStudy (shane_b):

He was just converting it to military time...so 4pm is 1600...8am = 0800...etc. It makes it slightly easier to figure out the hours between two times.

OpenStudy (anonymous):

Thanks